OpenStudy (anonymous):

what is the square root of 400? what is the square root of 400? @Mathematics

OpenStudy (anonymous):

\[\sqrt{400}=\sqrt{4\cdot100}=\sqrt{4}\cdot\sqrt{100}\]

OpenStudy (anonymous):

obviously 20

OpenStudy (anonymous):

20

OpenStudy (anonymous):

20